OpenStudy (anonymous):

The sum of twice a number and 18 less the number is the same as the difference between -34 and the number. What is the number? I wrote it like this; x+18=-34-x x+18-18=-34-18-x x=-52-x x+x=-52-x+x 2x=-52 2/2x = -52/2 x = -26 I have translated it incorrectly but I am unsure where.

OpenStudy (anonymous):

Is it supposed to have been: 2x+x-18= -34-x 3x-18=-34-x 3x+x-18= -34-x+x 4x-18=-34 4x-18+18= -34+18 4x= -16 4/4x= -16/4 x=-4 ????????????